OpenStudy (anonymous):

Consider a quadratic equation 2x^(2)-mx+(m+1)=0. If the roots of the equations are alpha and 2(alpha)+1, find the possible value(s) of m.

OpenStudy (anonymous):

\(\alpha \) and \(2\alpha +1\)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

for the equation: ax^2+bx+c=0 sum of roots r_1+r_2=-b/a product of roots: r_1*r_2=c/a

OpenStudy (anonymous):

r_1+r_2???

OpenStudy (anonymous):

sum of roots is \(-\frac{b}{a}\) in your case \[\alpha +2\alpha +1=\frac{m}{2}\] product of the roots is \(\frac{c}{a}\) you get \(\alpha (2\alpha +1)=\frac{m+1}{2}\)

OpenStudy (anonymous):

r_1 is 1º root, r_2 2º root

OpenStudy (anonymous):

by which i mean (and @myko means) the sum of the roots of \[ax^2+bx+c=0\]

OpenStudy (anonymous):

you get myko is typing a reply… Type your reply \[3\alpha +1=\frac{m}{2}\] or \[m=6\alpha +2\] \[\alpha^2+\alpha =\frac{m+1}{2}\] or \[2\alpha ^2+2\alpha -1=m\]

OpenStudy (anonymous):

set them equal and solve for \(m\) in terms of \(\alpha\)

OpenStudy (anonymous):

\[3\alpha+1=m\] \[\alpha=\frac{ m-1 }{ 3 }\] ---(1) \[2\alpha ^{2}+\alpha=m+1\]---(2) Sub. (1) into (2),we have